# A Two-Stage Pretraining-Finetuning Framework for Treatment Effect Estimation with Unmeasured Confounding

### v1

@@ 6. Conclusion
This paper studies the CATE estimation in the presence of unmeasured confounding fusing both large-scale OBS data and small-scale RCT data. We propose a two-stage pretraining-finetuning framework to tackle the overfitting problem caused by the small-scale RCT data. Specifically, the foundational representation learned in the first stage is used to adjust for the measured confounding in the OBS data. The augmented representation learned in the second stage is used to adjust for the unmeasured confounding guided by the RCT data. To avoid overfitting caused by the small-scale RCT data in the second stage, instead of training a separate network, we further propose to partially initialize the network parameters from the pretrained network from the first stage. Compared to the previous CATE estimation methods that combine OBS and RCT data, our approach has the advantage of not restricting the data-generating process (e.g., linearity or additive noise assumptions) and not suffering from overfitting. The extensive semi-synthetic and real-world experiments conducted on two widely-used public datasets demonstrate the superiority of our method.

### v2

@@ 5. Conclusion
This paper studies the CATE estimation problem in the presence of hidden confounding for fusing large-scale OBS data and small-scale RCT data. We propose a two-stage pretraining-finetuning framework to tackle the overfitting problem caused by the small-scale RCT data. Specifically, the foundational representation learned in the first stage is used to adjust for the measured confounding bias in the OBS data. The augmented representation learned in the second stage is used to mitigate for the hidden confounding bias guided by the RCT data. To avoid overfitting caused by the small-scale RCT data in the second stage, instead of training a separate network, we propose to partially initialize the network parameters from the pretrained network from the first stage. Compared to the previous CATE estimation methods that combine OBS and RCT data, our approach has the advantage of not restricting the data-generating process (e.g., linearity or additive noise assumptions) and suffering from overfitting. The experiments conducted in the real-world datasets demonstrate the superiority of our approach.